OpenStudy (anonymous):

What is the Quotient (x3+7x2+5x-20)÷(x+3)

OpenStudy (anonymous):

\[ \frac{x^3+7 x^2+5 x-20}{x+3}= x^2+4 x -7 +\frac{1}{x+3} \] So the quotient is \[ x^2+4 x -7 \]